# Shuffle Products, Degenerate Affine Hecke Algebras, and Quantum Toda Lattice

### v1

@@ Introduction
Shuffle algebras are ubiquitous in representation theory. To name a few applications, they are related to infinite-dimensional quantum groups (including quantum toroidal algebras), Hall algebras, canonical bases, and so on. They were introduced by Feigin-Odesskii in a series of papers [FeiginOdesskiiSklyanin, OdesskiiFeiginEllipticDeformation, FeiginOdesskii]. In the simplest case, a shuffle algebra is, as a vector space, the direct sum $\bigoplus_{k}\mathcal{O}(Y_{1},\ldots,Y_{k})^{S_{k}}$ of suitably defined symmetric functions in $k$ variables, with the product
@@
In this paper, we study the rational case. The elliptic one above admits a degeneration to it [EndelmanHodgesTwistedSU] so that
@@
The first goal of the paper is to relate these shuffle algebras to degenerate affine Hecke algebras whose definition we recall in Section 2.2. The idea is as follows: observe that without the theta factor, the shuffle product is simply the multiplication on the symmetric algebra of the vector space $\mathbf{C}[Y_{1}]$ or its $N$-th truncation. Instead of deforming it via inserting a factor, we can deform the symmetrization itself, that is, the action of the symmetric group. One such deformation is provided by the Demazure operators, so that a simple transposition acts on $\mathbf{C}[Y_{1},\ldots,Y_{k}]$ by
@@
The main result of the first part of the paper is the following, see Theorem 2.11.
@@
We must admit that, strictly speaking, the result is by no means new: for instance, it can considered as simply a reformulation of the known relation between explicit formulas for Hall-Littlewood polynomials and affine Hecke algebras [Macdonald]. However, what we find interesting is that it is possible to treat both rational and trigonometric cases simultaneously (related respectively to degenerate and non-degenerate affine Hecke algebras) via the so-called Shibukawa-Ueno $R$-operators whose definition we recall in Section 2.1. Those are particular solutions of the quantum Yang-Baxter equation
@@
introduced in [ShibukawaUeno] acting on the space of suitably defined functions, and they come in three types depending on one of the factors (1). At the same time, there is a canonical affinization procedure [KSQuantumGroups] that produces a solution to the quantum Yang-Baxter equation out of a representation of a symmetric group (or a finite Hecke algebra). It is an easy calculation that the affinization of the Demazure action above is essentially the rational Shibukawa-Ueno $R$-operator (and, similarly, the affinization of the action of the finite Hecke algebra on the spherical module of the affine Hecke algebra is its trigonometric version, although we do not calculate it in the paper and leave it as remark). Then the explicit formula for the symmetrization of the Demazure action can be computed via a certain combination of the $R$-operators which is done in Proposition 2.3.
@@
Interestingly, as we mentioned, there is an elliptic version of the Shibukawa-Ueno operator. Unfortunately, we do not know if and how it might be related to elliptic affine Hecke algebras [GinzburgKapranovVasserot]. Curiously, the elliptic case admits a dual description on the $\mathrm{GL}_{N}$-side, in the sense of Fourier-Mukai transform on an elliptic curve. Namely, by [FelderPasquier], the elliptic Shibukawa-Ueno $R$-operator is gauge-equivalent to Belavin’s $R$-matrix [Belavin, CherednikRmatrix] whose quasi-periodic properties can be interpreted via an essentially unique stable vector bundle of rank $N$ and degree 1 on an elliptic curve. A natural question is: what is the meaning of the shuffle product on the $\mathrm{GL}_{N}$-side, at least in the rational case?
@@
This question motivates the second part of the paper where we apply the construction of the first part to the quantum Toda lattice for $\mathrm{GL}_{N}$. The idea to use degenerate affine Hecke algebras is inspired by the results of [CautisKamnitzer] where the authors constructed and proved a K-theoretic version of the geometric Satake equivalence for a certain subcategory of the $q$-analog of Harish-Chandra bimodules (we recall the rational versions below). In the Kazhdan-Kostant formulation [Kostant], consider a non-degenerate character $\psi$ of the Lie algebra $\mathfrak{n}_{+}$ of the subgroup of upper unitriangular matrices $N_{+}$. The action of $\mathfrak{n}_{+}$ via right- and left-invariant vector fields gives rise to a homomorphism $\mathrm{U}(\mathfrak{n}_{+})\otimes\mathrm{U}(\mathfrak{n}_{+})\rightarrow\mathrm{D}(\mathrm{GL}_{N})$ into the algebra of differential operators. Then the Toda Hamiltonians are realized as the image of bi-invariant differential operators, which can be identified with the center $\mathbf{Z}_{\mathfrak{gl}_{N}}$ of $\mathrm{U}(\mathfrak{gl}_{N})$, inside the quantum Hamiltonian reduction $N^{L}_{+}\,\backslash\!\!\backslash_{-\psi}\,\mathrm{D}(\mathrm{GL}_{N})/\!\!/_{\psi}\,N_{+}^{R}$ with respect to the latter embedding and the character $(-\psi,\psi)$. This is an example of the Kostant-Whittaker reduction for Harish-Chandra bimodules studied in [BezrukavnikovFinkelberg].
@@
According to [BravermanFinkelbergNakajima, Appendix B], the Toda lattice can be identified with a certain quotient $\mathrm{Y}^{0}_{-2N}(\mathfrak{sl}_{2})$ of the shifted Yangian $\mathrm{Y}_{-2N}(\mathfrak{sl}_{2})$ whose definition we recall in Section 2.3. For introduction, we will only use the fact that it is generated by the coefficients of the series
@@
and that the coefficients of $x_{+}(u)$ generate the subalgebra isomorphic to the shuffle algebra $\mathcal{S}^{\mathrm{rat}}$ [TsymbaliukPBWD]. Therefore, it gives a $\mathrm{GL}_{N}$-presentation of $\mathcal{S}^{\mathrm{rat}}$. However, citing [FKPRW], “we must admit that the identification of the quantum Toda [lattice] for $\mathrm{GL}_{N}$ and the [truncated] shifted Yangian for $\mathfrak{sl}_{2}$ (purely algebraic objects) goes through a topological medium: equivariant homology of the affine Grassmannian of $\mathrm{GL}_{N}$”; this additional step is the derived Satake equivalence of [BezrukavnikovFinkelberg]. For instance, while $\mathrm{Y}_{-2N}(\mathfrak{sl}_{2})$ is defined via generators and relations, this identification does not specify explicitly the images of the former.
@@
as well as related constructions, in a purely algebraic and explicit way, bypassing the topological medium, and “refine” it to a mirabolic version of the Kostant-Whittaker reduction. It is based on another classical fact. For any $\mathrm{GL}_{N}$-representation $V$, the free left $\mathrm{U}(\mathfrak{gl}_{N})$-module $\mathrm{U}(\mathfrak{gl}_{N})\otimes V$ has a right $\mathrm{U}(\mathfrak{gl}_{N})$ action defined from the diagonal group action of $\mathrm{GL}_{N}$ on the tensor product. This is an example of a (free) Harish-Chandra bimodule whose definition and properties we review in Section 3. For the tensor product of the vector representation $V=(\mathbf{C}^{N})^{\otimes k}$, there is an action of the degenerate affine Hecke algebra $\mathfrak{H}_{k}$ on $\mathrm{U}(\mathfrak{gl}_{N})\otimes(\mathbf{C}^{N})^{\otimes k}$ respecting the $\mathrm{U}(\mathfrak{gl}_{N})$-bimodule structure. In particular, its spherical subalgebra $\mathcal{S}_{k}\cong e_{k}\mathfrak{H}_{k}e_{k}$ acts on the $S_{k}$-invariants that can be identified with the free Harish-Chandra bimodule $\mathrm{U}(\mathfrak{gl}_{N})\otimes S^{k}\mathbf{C}^{N}$ associated to the $k$-th symmetric power. By restricting the action on the highest-weight vector for all $k$, one obtains a map of vector spaces
@@
where the target is the quantum Hamiltonian reduction of the semi-direct product with $\mathfrak{gl}_{N}$ acting on the symmetric algebra $S^{\bullet}\mathbf{C}^{N}$ in the standard way. In particular, it is an algebra as well. If we denote by $\mathfrak{m}_{N}$ the mirabolic subalgebra of the subgroup of $\mathrm{GL}_{N}$ preserving the dual of the last basis vector in $\mathbf{C}^{N}$ (see Section 5.3 for precise definition), then it turns out that this map restricts to a similarly defined quantum Hamiltonian reduction
@@
The following is Theorem 5.14 and Theorem 5.9; here $\mathrm{Y}_{-2N}(\mathfrak{sl}_{2})^{\geq 0}$ is the non-negative part of the shifted Yangian. In fact, the explicit formulas for the second homomorphism are close to another presentation of the quantization of Zastava spaces for $\mathfrak{sl}_{2}$ from [FinkelbergRybnikov, Section 3.22].
@@
Unfortunately, we do not know if and how this result can be related to the aforementioned stable bundles on an elliptic curve. We expect (or, rather, hope) that it could viewed as a quantum rational version of a certain algebra of Hecke modifications of semistable bundles of degree 0, see Remark 5.15. In general, it would be interesting to interpret this result at least on the classical level in terms of elliptic analog of the Kostant slice [FriedmanMorgan].
Besides degenerate affine Hecke algebras, the other main tool to prove the theorem is the Kirillov projector $P_{\psi}$ introduced in [KalmykovYangians]. This is an element in a certain completion of $\mathrm{U}(\mathfrak{m}_{N})$ that acts on Whittaker modules and projects onto the space of Whittaker vectors (that is, on which $\mathfrak{n}_{+}$ acts with character $\psi$); as such, it is an analog of the classical extremal projector [AsherovaSmirnovTolstoy] in the case of trivial character. For instance, the Kirillov projector gives a canonical isomorphism $V\xrightarrow{\sim}\mathrm{U}(\mathfrak{m}_{N})\otimes V/\!\!/_{\psi}\,N_{+}$ of any $\mathrm{GL}_{N}$-representation $V$ with the mirabolic version of the Kostant-Whittaker reduction. It turns out that this isomorphism has rather peculiar properties with respect to the degenerate affine Hecke algebra action mentioned above, see Section 5.3. For example, under the isomorphism (3), the standard monomial basis of $S^{\bullet}\mathbf{C}^{N}$ is mapped to a rational version of the Hall-Littlewood polynomials, see Example 5.12.
We extend this homomorphism to the Toda lattice. Naturally, $\mathrm{GL}_{N}$ is embedded into the matrix space; denote by $X=(x_{ji})$ the (transposed) matrix with $x_{ij}\in\mathcal{O}(\mathrm{GL}_{N})$ the $(i,j)$-matrix entry function. Its inverse $X^{-1}$ is well-defined in $\mathcal{O}(\mathrm{GL}_{N})$. Also, denote by $E^{R}=(E_{ij}^{R})$ the matrix composed of left-invariant vector fields associated to the standard generators of the Lie algebra $\mathfrak{gl}_{N}$. The following is Theorem 7.9. It would interesting to compare this explicit form with the approach of [FKPRW] via coproducts on both sides.
@@
One of the main ideas of the proof is to interpret the GKLO homomorphism from $\mathrm{Y}_{-2N}(\mathfrak{sl}_{2})$ into a certain localization ${}^{N}\tilde{\mathcal{A}}$ of the algebra of difference operators, introduced in [GKLO] for non-shifted Yangians and later studied in [KWWY, BravermanFinkelbergNakajima] in general case, via a finite analog of the Miura transform [GinzburgKazhdan], see Section 6. In general, for any Harish-Chandra bimodule $X$, it relates the Kostant-Whittaker reduction $X/\!\!/_{\psi}\,N_{+}$ to its version with the trivial character, the parabolic restriction $X^{\mathrm{gen}}/\!\!/N_{-}$, which we review in Section 4 (here, the superscript “gen” corresponds to a certain localization), akin to the Harish-Chandra homomorphism $\mathbf{Z}_{\mathfrak{gl}_{N}}\rightarrow\mathrm{U}(\mathfrak{h}_{N})\cong(\mathfrak{n}_{-}\backslash\mathrm{U}(\mathfrak{gl}_{N}))^{N_{-}}$. In particular, it gives an injective homomorphism
@@
The following is Theorem 7.7 and the second part of Theorem 7.9
@@
In Section 7.2, we apply this theorem to get simplest cases of the monopole operators, which are rational versions of certain degenerations of Macdonald difference operators, as the images of particular minor functions on $\mathrm{GL}_{N}$. We express the latter in terms of the shuffle algebra similarly to the geometric and $q$-setting of [FinkelbergTsymbaliuk] and [TsymbaliukGKLO]. Finally, in Section 7.1, we show that this isomorphism coincides with the one in [BravermanFinkelbergNakajima, Appendix B] under the derived Satake isomorphism of [BezrukavnikovFinkelberg] between $N^{L}_{+}\,\backslash\!\!\backslash_{-\psi}\,\mathrm{D}(\mathrm{GL}_{N})/\!\!/_{\psi}\,N_{+}^{R}$ and the equivariant Borel-Moore homology $H^{\mathrm{GL}_{N}[[t]]\rtimes\mathbf{C}^{*}}_{\bullet}(\mathrm{Gr})$ of the affine Grassmannian of $\mathrm{GL}_{N}$.
We would like to mention that we actively use the formulas (adapted to setting of the paper) from the paper [MolevTransvectorAlgebras] that studied Olshanski centralizer construction from the point of view of Micklesson algebras whose particular case we recall in Section 4. It would be interesting to find a direct conceptual link between two subjects.
@@
In Section 1, we introduce the necessary notations concerning $\mathrm{GL}_{N}$ and differential operators $\mathrm{D}(\mathrm{GL}_{N})$ as well quantum minors. In Section 2, we recall basic properties of shifted Yangians, degenerate affine Hecke algebras, and relate the latter to Shibukawa-Ueno $R$-operators and shuffle algebras. In Section 3, we recall the definition of Harish-Chandra bimodules and degenerate affine Hecke algebras action on the free ones associated to tensor powers of the vector representation. In Section 4, we study quantum Hamiltonian reduction with a trivial character, the parabolic restriction. In Section 5, we study the Kostant-Whittaker reduction and its mirabolic version relating them to degenerate affine Hecke algebras, shuffle algebras, and (the positive part of) shifted Yangians. In Section 6, we relate two reductions via finite Miura transform and interpret it as a GKLO map. In Section 7, we prove an isomorphism between truncated shifted Yangians and Toda lattice, compare it with the geometric identification via homology of the affine Grassmannian, and interpret monopole operators in terms certain minor functions. In Appendix A, we prove some technical lemmas from the main sections. While in the main body of the paper we deal with the vector representation, in Appendix B, we present the corresponding facts for the dual that are needed for the full Toda lattice.

### v2

@@ Introduction
Shuffle algebras are ubiquitous in representation theory. To name a few applications, they are related to infinite-dimensional quantum groups (including quantum toroidal algebras), Hall algebras, canonical bases, and so on. They were introduced by Feigin-Odesskii in a series of papers [OF89, OF97, FO98]. In the simplest case, a shuffle algebra is, as a vector space, the direct sum $\bigoplus_{k}\mathcal{O}(Y_{1},\ldots,Y_{k})^{S_{k}}$ of suitably defined symmetric functions in $k$ variables, with the product
@@
In this paper, we study the rational case. The elliptic one above admits a degeneration to it [EH04] so that
@@
The first goal of the paper is to relate these shuffle algebras to degenerate affine Hecke algebras whose definition we recall in Section˜2.2. The idea is as follows: observe that without the theta factor, the shuffle product is simply the multiplication on the symmetric algebra of the vector space $\mathbf{C}[Y_{1}]$ or its $N$-th truncation. Instead of deforming it via inserting a factor, we can deform the symmetrization itself, that is, the action of the symmetric group. One such deformation is provided by the Demazure operators, so that a simple transposition acts on $\mathbf{C}[Y_{1},\ldots,Y_{k}]$ by
@@
The main result of the first part of the paper is Theorem˜2.11.
@@
We must admit that the result is by no means new: for instance, its trigonometric version can considered as simply a reformulation of the known relation between explicit formulas for Hall-Littlewood polynomials and affine Hecke algebras [Mac95]. Our main contribution is an alternative proof via the Shibukawa-Ueno $R$-operators whose definition we recall in Section˜2.1. Those are particular solutions of the quantum Yang-Baxter equation
@@
introduced in [SU92] that act on the space of suitably defined functions and take as an input any solution to the three-term relation (2.1) (in particular, one of the functions (1)). At the same time, there is a canonical affinization procedure [KS97] that produces a solution to the quantum Yang-Baxter equation out of a representation of a symmetric group (or a finite Hecke algebra). It is an easy calculation that the affinization of the Demazure action above is essentially the rational Shibukawa-Ueno $R$-operator (and, similarly, the affinization of the action of the finite Hecke algebra on the spherical module of the affine Hecke algebra is its trigonometric version, although we do not calculate it in the paper and leave it as remark). Then the explicit formula for the symmetrization of the Demazure action can be computed via a certain combination of the $R$-operators which is done in Proposition˜2.3.
@@
Interestingly, as we mentioned, there is an elliptic version of the Shibukawa-Ueno operator. Unfortunately, we do not know if and how it might be related to elliptic affine Hecke algebras [GKV97]. Curiously, the elliptic case admits a dual description on the $\mathrm{GL}_{N}$-side, in the sense of Fourier-Mukai transform on an elliptic curve. Namely, by [FP94], the elliptic Shibukawa-Ueno $R$-operator is gauge-equivalent to Belavin’s $R$-matrix [Bel80, Che86] whose quasi-periodic properties can be interpreted via an essentially unique stable vector bundle of rank $N$ and degree 1 on an elliptic curve. A natural question is: what is the meaning of the shuffle product on the $\mathrm{GL}_{N}$-side?
@@
We approach this question for the rational case in the second part of the paper where we apply the construction of the first part to the quantum Toda lattice for $\mathrm{GL}_{N}$. The idea to use degenerate affine Hecke algebras is inspired by the results of [CK18] where the authors constructed and proved a K-theoretic version of the geometric Satake equivalence for a certain subcategory of the $q$-analog of Harish-Chandra bimodules (we recall the rational versions below). Following the Kazhdan-Kostant formulation of this integrable system [Kos78], let us consider a non-degenerate character $\psi$ of the Lie algebra $\mathfrak{n}_{+}$ of the subgroup of upper unitriangular matrices $N_{+}$. The action of $\mathfrak{n}_{+}$ via right- and left-invariant vector fields gives rise to a homomorphism $\mathrm{U}(\mathfrak{n}_{+})\otimes\mathrm{U}(\mathfrak{n}_{+})\rightarrow\mathrm{D}(\mathrm{GL}_{N})$ into the algebra of differential operators. Then the Toda Hamiltonians are realized as the image of bi-invariant differential operators, which can be identified with the center $\mathbf{Z}_{\mathfrak{gl}_{N}}$ of $\mathrm{U}(\mathfrak{gl}_{N})$, inside the quantum Hamiltonian reduction $N^{L}_{+}\,\backslash\!\!\backslash_{-\psi}\,\mathrm{D}(\mathrm{GL}_{N})/\!\!/_{\psi}\,N_{+}^{R}$ with respect to the latter embedding and the character $(-\psi,\psi)$. This is an example of the Kostant-Whittaker reduction for Harish-Chandra bimodules studied in [BF08].
@@
According to [BFN19, Appendix B], the Toda lattice can be identified with a certain quotient $\mathrm{Y}^{0}_{-2N}(\mathfrak{sl}_{2})$ of the shifted Yangian $\mathrm{Y}_{-2N}(\mathfrak{sl}_{2})$ whose definition we recall in Section˜2.3. For introduction, we will only use the fact that it is generated by the coefficients of the series
@@
and that the coefficients of $x_{+}(u)$ generate the subalgebra isomorphic to the shuffle algebra $\mathcal{S}^{\mathrm{rat}}$ [Tsy21]. Therefore, it gives a $\mathrm{GL}_{N}$-presentation of $\mathcal{S}^{\mathrm{rat}}$. However, citing [Fin+18], “we must admit that the identification of the quantum Toda [lattice] for $\mathrm{GL}_{N}$ and the [truncated] shifted Yangian for $\mathfrak{sl}_{2}$ (purely algebraic objects) goes through a topological medium: equivariant homology of the affine Grassmannian of $\mathrm{GL}_{N}$”; this additional step is the derived Satake equivalence of [BF08]. For instance, while $\mathrm{Y}_{-2N}(\mathfrak{sl}_{2})$ is defined via generators and relations, this identification does not specify explicitly the images of the former.
@@
as well as related constructions, in a purely algebraic and explicit way, bypassing the topological medium, and “refine” it to a mirabolic version of the Kostant-Whittaker reduction. It is based on another classical fact. For any $\mathrm{GL}_{N}$-representation $V$, the free left $\mathrm{U}(\mathfrak{gl}_{N})$-module $\mathrm{U}(\mathfrak{gl}_{N})\otimes V$ has a right $\mathrm{U}(\mathfrak{gl}_{N})$ action defined from the diagonal group action of $\mathrm{GL}_{N}$ on the tensor product. This is an example of a (free) Harish-Chandra bimodule whose definition and properties we review in Section˜3. For the tensor product of the vector representation $V=(\mathbf{C}^{N})^{\otimes k}$, there is an action of the degenerate affine Hecke algebra $\mathfrak{H}_{k}$ on $\mathrm{U}(\mathfrak{gl}_{N})\otimes(\mathbf{C}^{N})^{\otimes k}$ respecting the $\mathrm{U}(\mathfrak{gl}_{N})$-bimodule structure. In particular, its spherical subalgebra $\mathcal{S}_{k}\cong e_{k}\mathfrak{H}_{k}e_{k}$ acts on the $S_{k}$-invariants that can be identified with the free Harish-Chandra bimodule $\mathrm{U}(\mathfrak{gl}_{N})\otimes S^{k}\mathbf{C}^{N}$ associated to the $k$-th symmetric power. By restricting the action on the highest-weight vector for all $k$, one obtains a map of vector spaces
@@
where the target is the quantum Hamiltonian reduction of the semi-direct product with $\mathfrak{gl}_{N}$ acting on the symmetric algebra $S^{\bullet}\mathbf{C}^{N}$ in the standard way. In particular, it is an algebra as well. If we denote by $\mathfrak{m}_{N}$ the mirabolic subalgebra of the subgroup of $\mathrm{GL}_{N}$ preserving the dual of the last basis vector in $\mathbf{C}^{N}$ (see Section˜5.3 for precise definition), then it turns out that this map restricts to a similarly defined quantum Hamiltonian reduction
@@
The following is Theorem˜5.14 and Theorem˜5.9; here $\mathrm{Y}_{-2N}(\mathfrak{sl}_{2})^{\geq 0}$ is the non-negative part of the shifted Yangian. In fact, the explicit formulas for the second homomorphism are close to another presentation of the quantization of Zastava spaces for $\mathfrak{sl}_{2}$ from [FR14, Section 3.22].
@@
Unfortunately, we do not know if and how this result can be related to the aforementioned stable bundles on an elliptic curve. We expect (or, rather, hope) that it could viewed as a quantum rational version of a certain algebra of Hecke modifications of semistable bundles of degree 0, see Remark˜5.15. In general, it would be interesting to interpret this result at least on the classical level in terms of elliptic analog of the Kostant slice [FM00].
Besides degenerate affine Hecke algebras, the other main tool to prove the theorem is the Kirillov projector $P_{\psi}$ introduced in [Kal25]. This is an element in a certain completion of $\mathrm{U}(\mathfrak{m}_{N})$ that acts on Whittaker modules and projects onto the space of Whittaker vectors (that is, on which $\mathfrak{n}_{+}$ acts with character $\psi$); as such, it is an analog of the classical extremal projector [AST71] in the case of trivial character. For instance, the Kirillov projector gives a canonical isomorphism $V\xrightarrow{\sim}\mathrm{U}(\mathfrak{m}_{N})\otimes V/\!\!/_{\psi}\,N_{+}$ of any $\mathrm{GL}_{N}$-representation $V$ with the mirabolic version of the Kostant-Whittaker reduction. It turns out that this isomorphism has rather peculiar properties with respect to the degenerate affine Hecke algebra action mentioned above, see Section˜5.3. For example, under the isomorphism (3), the standard monomial basis of $S^{\bullet}\mathbf{C}^{N}$ is mapped to a rational version of the Hall-Littlewood polynomials, see Example˜5.12.
We extend this homomorphism to the Toda lattice. Naturally, $\mathrm{GL}_{N}$ is embedded into the matrix space; denote by $X=(x_{ji})$ the (transposed) matrix with $x_{ij}\in\mathcal{O}(\mathrm{GL}_{N})$ the $(i,j)$-matrix entry function. Its inverse $X^{-1}$ is well-defined in $\mathcal{O}(\mathrm{GL}_{N})$. Also, denote by $E^{R}=(E_{ij}^{R})$ the matrix composed of left-invariant vector fields associated to the standard generators of the Lie algebra $\mathfrak{gl}_{N}$. The following is Theorem˜7.9. It would interesting to compare this explicit form with the approach of [Fin+18] via coproducts on both sides.
@@
[…]
@@
The following is Theorem˜7.7 and the second part of Theorem˜7.9
@@
In Section˜7.2, we apply this theorem to get simplest cases of the monopole operators, which are rational versions of certain degenerations of Macdonald difference operators, as the images of particular minor functions on $\mathrm{GL}_{N}$. We express the latter in terms of the shuffle algebra similarly to the geometric and $q$-setting of [FT19] and [Tsy23]. Finally, in Section˜7.1, we show that this isomorphism coincides with the one in [BFN19, Appendix B] under the derived Satake isomorphism of [BF08] between $N^{L}_{+}\,\backslash\!\!\backslash_{-\psi}\,\mathrm{D}(\mathrm{GL}_{N})/\!\!/_{\psi}\,N_{+}^{R}$ and the equivariant Borel-Moore homology $H^{\mathrm{GL}_{N}[[t]]\rtimes\mathbf{C}^{*}}_{\bullet}(\mathrm{Gr})$ of the affine Grassmannian of $\mathrm{GL}_{N}$.
We would like to mention that we actively use the formulas (adapted to setting of the paper) from the paper [Mol02] that studied Olshanski centralizer construction from the point of view of Micklesson algebras whose particular case we recall in Section˜4. It would be interesting to find a direct conceptual link between two subjects.
@@
In Section˜1, we introduce the necessary notations concerning $\mathrm{GL}_{N}$ and differential operators $\mathrm{D}(\mathrm{GL}_{N})$ as well quantum minors. In Section˜2, we recall basic properties of shifted Yangians, degenerate affine Hecke algebras, and relate the latter to Shibukawa-Ueno $R$-operators and shuffle algebras. In Section˜3, we recall the definition of Harish-Chandra bimodules and degenerate affine Hecke algebras action on the free ones associated to tensor powers of the vector representation. In Section˜4, we study quantum Hamiltonian reduction with a trivial character, the parabolic restriction. In Section˜5, we study the Kostant-Whittaker reduction and its mirabolic version relating them to degenerate affine Hecke algebras, shuffle algebras, and (the positive part of) shifted Yangians. In Section˜6, we relate two reductions via finite Miura transform and interpret it as a GKLO map. In Section˜7, we prove an isomorphism between truncated shifted Yangians and Toda lattice, compare it with the geometric identification via homology of the affine Grassmannian, and interpret monopole operators in terms certain minor functions. In Appendix˜A, we prove some technical lemmas from the main sections. While in the main body of the paper we deal with the vector representation, in Appendix˜B, we present the corresponding facts for the dual that are needed for the full Toda lattice.